| <p>Frank has been tuning the beam since 08:00</p>
<p>detectors moved fully out</p>
<p>11:20</p>
<p>New run (R68) started at 11:20. Put into no storage mode as rates are high during beam tuning. Cut off the end of last nights run when analysing</p>
<p>System checks OK</p>
<p> </p>
<p>13.20 DSSSD bias & leakage current OK - attachment 1</p>
<p>FEE64 temperatures OK - attachment 2</p>
<p>ADC data item stats - attachment 3</p>
<p>per FEE64 1.8.W spectra - attachments 4-5</p>
<p>per FEE64 Rate spectra - attachment 6</p>
<p> aida16 producing HEC data - Chck ASIC load to fix</p>
<p>TapeServer - no storage mode</p>
<p> </p>
<p>13.35 aida01-aida04 c. 250k ADC data items/s - will await completion of current beam/target scan to fix by cycing Left DSSSD motion</p>
<p> </p>
<p>14.00 CB cycles Left DSSSD motion to fix aida01-aida04 rate issue</p>
<p> </p>
<p>CRYRING vacuum & hall temprature - attachments 7-8</p>
<p> </p>
<p>14.28 New run data file G22-201/R69</p>
<p>Beam/target overlap optimised - details to follow</p> |